OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! A twelve-gallon gas tank is thirty percent full. Which proportion could be used to determine how many gallons of gasoline are in the tank? Select one: A. twelve over thirty equals x over one hundred B. thirty over x equals twelve over one hundred C. x over twelve equals thirty over one hundred D. x over one hundred equals thirty over twelve

OpenStudy (anonymous):

@isaac4321

OpenStudy (anonymous):

Ooh I Think I know this one

OpenStudy (anonymous):

I believe it's c

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

Yea lets say that

OpenStudy (anonymous):

Yes I believe it is c also.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm 51 percent sure

OpenStudy (anonymous):

Let x= the number of gallons in the tank. Since there are 12 gallons the amount of gas left is(30/100)*12 Knowing that (30/100)*12=x You can divide both sides of the equation by 12 and get. C) x/12= 30/100. I found this answer on line I hope it helps.

OpenStudy (anonymous):

So it's c

OpenStudy (anonymous):

Yes it's C.

OpenStudy (anonymous):

Sorry guys my bad